# Evaluating Generative Ad Hoc Information Retrieval

### v2

@@ 3.4.3. Accumulation Model for Generative IR
To combine gain and discount values over all considered statements, we argue in favor of the accumulation model of expected total utility (Carterette, 2011; Moffat et al., 2013). It considers the total utility a searcher accumulates from the whole response. Alternatively, measures could be based around estimating the total ‘cost’ of accruing information from the response in terms of the effort expended (Carterette, 2011). However, we argue that, since this effort is comparatively small in text SERPs, optimizing for it is not suitable for reliably differentiating systems in evaluation.
@@ 4. Operationalizing Evaluation
This section considers operationalizations of the proposed user model. The goal is to take stake in what possibilities exist for each step of the process, in an effort to illustrate the required components and how they can be implemented. These considerations are summarized in Figure 6, with each component (Figure rows) described in a subsection below. The experimental setup encompasses a document collection, a set of topics reflecting the search task, and a set of generative retrieval systems to be evaluated (Section 4.1). Their responses to queries are (optionally) split into statements using a segmentation approach (Section 4.2). Statements are then assessed for their utility, distinguishing between assessment without prior reference, and assessment in relation to prior reference material (Section 4.3). Given annotations and an evaluation measure, the systems can then be ranked with respect to their effectiveness as indicated by an aggregated score (Section 4.4). In each of these four steps, we survey relevant literature and juxtapose proposed evaluation processes with regard to their advantages and disadvantages in the context of the assumed user model.
@@ Acknowledgements.
This publication has been partially supported by the ScaDS.AI Center for Scalable Data Analytics and Artificial Intelligence, funded by the Federal Ministry of Education and Research of Germany and by the Sächsische Staatsministerium für Wissenschaft, Kultur und Tourismus, by a Research Fellowship for Harrisen Scells from the Alexander von Humboldt Foundation, and by the OpenWebSearch.eu project, funded by the European Union (GA 101070014).References
@@
- Manning et al. (2008) Christopher D. Manning, Prabhakar Raghavan, and Hinrich Schütze. 2008. Introduction to Information Retrieval. Cambridge University Press.

### v3

@@ 3.4.3. An Accumulation Model for Generative IR
To combine gain and discount values over the statements of a response, we argue in favor of expected total utility accumulation (Carterette, 2011; Moffat et al., 2013). It considers the total utility a searcher accumulates from the whole response. Alternatively, measures could be based around estimating the total ‘cost’ of accruing information from the response in terms of the effort expended (Carterette, 2011). However, we argue that the effort is comparatively small in text SERPs so that optimizing for it is not that suitable to reliably differentiate systems in evaluation.
@@ 4. Operationalizing Evaluation
This section considers operationalizations of the proposed user model. The goal is to take stake in what possibilities exist for each step of the process, in an effort to illustrate the required components and how they can be implemented. These considerations are summarized in Figure 6, with each component (rows in the figure) described in a subsection below. The experimental setup encompasses a document collection, a set of topics reflecting the search task, and a set of generative retrieval systems to be evaluated (Section 4.1). Their responses to queries are (optionally) split into statements using a segmentation approach (Section 4.2). Statements are then assessed for their utility, distinguishing between assessment without prior reference, and assessment in relation to prior reference material (Section 4.3). Given annotations and an evaluation measure, the systems can then be ranked with respect to their effectiveness as indicated by an aggregated score (Section 4.4). In each of these four steps, we survey relevant literature and juxtapose proposed evaluation processes with regard to their advantages and disadvantages in the context of the assumed user model.
@@ Acknowledgements.
This publication has been partially supported by the ScaDS.AI Center for Scalable Data Analytics and Artificial Intelligence, funded by the Federal Ministry of Education and Research of Germany and by the Sächsische Staatsministerium für Wissenschaft, Kultur und Tourismus; by a Research Fellowship for Harrisen Scells from the Alexander von Humboldt Foundation; and by the OpenWebSearch.eu project, funded by the European Union (GA 101070014).References
@@
- Zerhoudi et al. (2022) Saber Zerhoudi, Sebastian Günther, Kim Plassmeier, Timo Borst, Christin Seifert, Matthias Hagen, and Michael Granitzer. 2022. The SimIIR 2.0 Framework: User Types, Markov Model-Based Interaction Simulation, and Advanced Query Generation. In Proceedings of the 31st ACM International Conference on Information & Knowledge Management, Atlanta, GA, USA, October 17–21, 2022, Mohammad Al Hasan and Li Xiong (Eds.). ACM, 4661–4666.